# Mendelian Randomization Methods for Causal Inference: Estimands, Identification and Inference

### v1

@@ 4.2 ALICE model is widely used in MR studies
where $\gamma_{j}$ represents the IV strength of $j$th genetic instrument. Note that the error term $\delta_{i}$ in the exposure model (4) might be correlated with the error term $\varepsilon_{i}$ in the outcome model (3) due to unmeasured confounders. By plugging in the exposure model (4) into the outcome model (3), we can obtain the reduced-form model for the outcome (Small,, 2007; Guo et al.,, 2018):
@@ 4.3 Practical limitations and interpretational caveats of the ALICE model
Although the ALICE model provides a useful framework for causal inference with potentially invalid instrumental variables, it is subject to several important limitations and interpretational caveats. First, the causal effect defined in the ALICE model should be interpreted as a constant treatment effect of a one-unit increase in the exposure on the outcome (Holland,, 1986; Small,, 2007; Kang et al.,, 2016). However, the constant treatment effect assumption may not hold in real-world settings where treatment effects may vary across subjects (Angrist,, 2004; Powers et al.,, 2018; Künzel et al.,, 2019). Second, the ALICE model assumes linearity not only in the causal effect but also in the violation of core IV assumptions. This linearity assumption might be violated when the underlying relationships are nonlinear, such as in complex genetic architectures (Veitia et al.,, 2013; Guindo-Mart´ınez et al.,, 2021; Sun et al., 2023a, ). Therefore, when applying the ALICE model, it is essential to carefully assess the plausibility of its assumptions within the context of the study and to interpret the resulting estimates with appropriate caution.
@@ 8.2 MR methods using summary-level data
where $\widehat{\sigma}_{\Gamma,j}$ is the standard error of $\widehat{\Gamma}_{j}$. The IVW estimator upweights genetic instruments with stronger IV-exposure associations (i.e., larger $\widehat{\gamma}_{j}^{2}$) and more precise IV-outcome associations (i.e., smaller $\widehat{\sigma}_{\Gamma,j}^{2}$). In addition, $\widehat{\beta}_{\text{IVW}}$ consistently estimates $\beta$ when all instruments are valid and mutually independent. When some genetic instruments violate the core IV assumptions, the IVW estimator is no longer consistent. To address this, several summary-level data MR methods have been proposed to obtain robust causal effect estimates even in the violation of core IV assumptions (Bowden et al.,, 2015, 2016; Verbanck et al.,, 2018; Zhao et al., 2019b, ; Burgess et al.,, 2020; Xu et al.,, 2023; Yao et al.,, 2024). We provide a list of commonly used software implementations with links in Supplementary Section S2.
@@ 9.2 Weak IV biases in one-sample and two-sample MR estimations
where $\sigma_{\delta,\varepsilon}$ is the covariance between $\bm{\delta}^{(s)}$ and $\bm{\varepsilon}^{(s)}$, $\sigma_{\delta}^{2}$ is the variance of $\bm{\delta}^{(s)}$, and $\|\bm{\gamma}\|^{2}=\sum_{j=1}^{p}\gamma_{j}^{2}$. We also provide the derivation of this approximate bias in Supplementary Section S1. On the other hand, the approxmate bias of ordinary least square (OLS) estimator $\widehat{\beta}_{\text{OLS}}^{(s)}$ using sample $s$ is
@@
$\mathbb{E}\left(\widehat{\beta}_{\text{IVW}}^{(s)}\right)-\beta\approx\frac{\sigma_{\delta,\varepsilon}}{n^{(s)}}\times\frac{\sum_{j=1}^{p}1/(\widehat{\sigma}^{(s)}_{\Gamma,j})^{2}}{\sum_{j=1}^{p}(\gamma_{j}^{2}+(\sigma_{\gamma,j}^{(s})^{2})/(\widehat{\sigma}^{(s)}_{\Gamma,j})^{2}},$
@@ 11 Applications in Real Datasets
We compare the following methods for estimating the causal effect $\beta$: (1) two-stage least squares (2SLS) (Wooldridge,, 2016); (2) Two-Stage Hard Thresholding (TSHT) (Guo et al.,, 2018); (3) Confidence Interval method for Instrumental Variable (CIIV) (Windmeijer et al.,, 2021); (4) Some Invalid Some Valid IV Estimator (sisVIVE) (Kang et al.,, 2016); (5) MRSquare (Sun et al., 2023a, ); MR Mixed-Scale Treatment Effect Robust Identification (MR-MiSTERI) (Liu et al.,, 2023); and MR with MAny weak Genetic Interactions for Causality (MR-MAGIC) (Zhang et al.,, 2025). For sisVIVE, we choose the tuning parameter via 10-fold cross-validation. For MRSquare, we set the minimum number of valid instruments to be $6$. The results are summarized in Figure 6.
@@
The increasingly available large-scale multi-omics data (e.g., epigenomics, transcriptomics, proteomics, and metabolomics data) enable us to perform omics MR (xMR, firstly coined in Yao et al., (2024)) to detect putative causal omics biomarkers for complex traits and diseases, thereby uncovering the underlying causal mechanisms. For a detailed, step‑by‑step tutorial on implementing commonly used xMR methods, please refer to Yao and Liu, (2025).
In this section, we apply several two-sample MR methods to perform xMR analysis, aiming to identify putative causal plasma proteins associated with the risk of Alzheimer’s disease. For the exposure, we use UK Biobank Pharma Proteomics Project (UKB‑PPP) summary statistics on 1,463 plasma proteins measured in 54,306 individuals (Sun et al., 2023b, ). For the outcome, we use summary statistics from a meta-analysis of GWASs for clinically diagnosed AD and AD by proxy, comprising 455,258 samples in total (Jansen et al.,, 2019). Genetic instruments for each protein are selected by applying a Bonferroni‐corrected threshold of $p\text{-value}<3.40\times 10^{-11}$, followed by LD clumping at threshold $r^{2}<0.01$, as described in Sun et al., 2023b . We compare the following two-sample MR methods in this application: (1) inverse-variance weighted (IVW) method (Bowden et al.,, 2017); (2) MR-Egger regression (Bowden et al.,, 2015); (3) MR using the Robust Adjusted Profile Score (MR-RAPS) (Zhao et al.,, 2020); (4) MR Pleiotropy RESidual Sum and Outlier test (MR-PRESSO) (Verbanck et al.,, 2018); (5) the weighted median method (Bowden et al.,, 2016); (6) the mode-based estimation (Hartwig et al.,, 2017); (7) MRMix (Qi and Chatterjee,, 2019); (8) the contamination mixture method (Burgess et al.,, 2020); and (9) MR with valid IV Selection and Post-selection Inference (MR-SPI) (Yao et al.,, 2024).
Proteins identified to be significantly associated with Alzheimer’s disease after Bonferroni correction (Bland and Altman,, 1995) are summarized in Figure S1. In Figure S1(A), we report the number of significant plasma proteins detected by each method, which ranges from 0 (MR‑PRESSO) to 14 (MRMix). MR‑PRESSO detects no significant proteins, likely because the small number of candidate IVs per protein limits its power to perform the outlier test and detect invalid instruments. In Figure S1(B), we list the 11 plasma proteins identified by at least two methods. Notably, the seven proteins identified by MR‑SPI correspond to the top seven proteins ranked by the number of supporting methods. For a more detailed discussion of these seven proteins including gene ontology analysis and AlphaFold3‑based structural prediction (Abramson et al.,, 2024), please refer to Yao et al., (2024).
@@ 12.1 Binary and survival outcomes
In epidemiological research, the binary and survival outcomes are common, yet current MR methods predominantly focus on continuous outcomes. For one-sample MR with individual-level data, establishing identification conditions for MR analysis with binary or survival outcomes is essential, particularly in the presence of invalid instruments (Clarke and Windmeijer,, 2010; Deng et al.,, 2023; Liu et al.,, 2025). For two-sample MR with summary-level data, the current standard practice is to directly apply existing two-sample MR methods under the ALICE model framework for continuous outcomes to analyze summary statistics of binary or survival outcomes, while the interpretation of the causal effect estimate obtained by this direct application is unclear and requires further justification (Zhao et al., 2019b, ).

### v2

@@ Abstract
Keywords: Causal genomics; Causal inference; Instrumental variables; Mendelian randomization; Omics data; Unmeasured confounding; UK Biobank
@@ 4.2 ALICE model is widely used in MR studies
where $\gamma_{j}$ represents the IV strength of $j$th genetic instrument.
Note that the error term $\delta_{i}$ in the exposure model (4) might be correlated with the error term $\varepsilon_{i}$ in the outcome model (3) due to unmeasured confounders. By plugging in the exposure model (4) into the outcome model (3), we can obtain the reduced-form model for the outcome (Small,, 2007; Guo et al.,, 2018):
@@ 4.3 Practical limitations and interpretational caveats of the ALICE model
Although the ALICE model provides a useful framework for causal inference with potentially invalid instrumental variables, it is subject to several important limitations and interpretational caveats. First, the causal effect defined in the ALICE model should be interpreted as a constant treatment effect of a one-unit increase in the exposure on the outcome (Holland,, 1986; Small,, 2007; Kang et al.,, 2016). However, the constant treatment effect assumption may not hold in real-world settings where treatment effects may vary across subjects (Angrist,, 2004; Powers et al.,, 2018; Künzel et al.,, 2019). Second, the ALICE model assumes linearity not only in the causal effect but also in the violation of core IV assumptions. This linearity assumption might be violated when the underlying relationships are nonlinear, such as in complex genetic architectures (Veitia et al.,, 2013; Guindo-Martínez et al.,, 2021; Sun et al., 2023a, ). Therefore, when applying the ALICE model, it is essential to carefully assess the plausibility of its assumptions within the context of the study and to interpret the resulting estimates with appropriate caution.
@@ 8.2 MR methods using summary-level data
where $\widehat{\sigma}_{\Gamma,j}$ is the standard error of $\widehat{\Gamma}_{j}$. The IVW estimator upweights genetic instruments with stronger IV-exposure associations (i.e., larger $\widehat{\gamma}_{j}^{2}$) and more precise IV-outcome associations (i.e., smaller $\widehat{\sigma}_{\Gamma,j}^{2}$). In addition, $\widehat{\beta}_{\text{IVW}}$ consistently estimates $\beta$ when all instruments are valid and mutually independent. When some genetic instruments violate the core IV assumptions, the IVW estimator is no longer consistent. To address this, several summary-level data MR methods have been proposed to obtain robust causal effect estimates even in the violation of core IV assumptions (Bowden et al.,, 2015, 2016; Verbanck et al.,, 2018; Zhao et al., 2019b, ; Burgess et al.,, 2020; Xu et al.,, 2023; Yao et al.,, 2024). We provide a list of commonly used software implementations with links in Supplementary Materials.
@@ 9.2 Weak IV biases in one-sample and two-sample MR estimations
where $\sigma_{\delta,\varepsilon}$ is the covariance between $\bm{\delta}^{(s)}$ and $\bm{\varepsilon}^{(s)}$, $\sigma_{\delta}^{2}$ is the variance of $\bm{\delta}^{(s)}$, and $\|\bm{\gamma}\|^{2}=\sum_{j=1}^{p}\gamma_{j}^{2}$. We also provide the derivation of this approximate bias in Supplementary Materials. On the other hand, the approxmate bias of ordinary least square (OLS) estimator $\widehat{\beta}_{\text{OLS}}^{(s)}$ using sample $s$ is
@@
$\mathbb{E}\left(\widehat{\beta}_{\text{IVW}}^{(s)}\right)-\beta\approx\frac{\sigma_{\delta,\varepsilon}}{n^{(s)}}\times\frac{\sum_{j=1}^{p}1/(\widehat{\sigma}^{(s)}_{\Gamma,j})^{2}}{\sum_{j=1}^{p}(\gamma_{j}^{2}+(\sigma_{\gamma,j}^{(s)})^{2})/(\widehat{\sigma}^{(s)}_{\Gamma,j})^{2}},$
@@ 11 Applications in Real Datasets
We compare the following methods for estimating the causal effect $\beta$: (1) two-stage least squares (2SLS) (Wooldridge,, 2016); (2) Two-Stage Hard Thresholding (TSHT) (Guo et al.,, 2018); (3) Confidence Interval method for Instrumental Variable (CIIV) (Windmeijer et al.,, 2021); (4) Some Invalid Some Valid IV Estimator (sisVIVE) (Kang et al.,, 2016); (5) MRSquare (Sun et al., 2023a, ); MR Mixed-Scale Treatment Effect Robust Identification (MR-MiSTERI) (Liu et al.,, 2023); and MR with MAny weak Genetic Interactions for Causality (MR-MAGIC) (Zhang et al.,, 2025). For sisVIVE, we choose the tuning parameter via 10-fold cross-validation. For MRSquare, we set the minimum number of valid instruments to be $6$. The code for this application is provided in “DataExamples.R” within the Supplementary Files. The results are summarized in Figure 6.
@@
The increasingly available large-scale multi-omics data (e.g., epigenomics, transcriptomics, proteomics, and metabolomics data) enable us to perform omics MR (xMR, firstly coined in Yao et al., (2024)), a methodology within the causal genomics field, to detect putative causal omics biomarkers for complex traits and diseases, thereby uncovering the underlying causal mechanisms. For a detailed, step‑by‑step tutorial on implementing commonly used xMR methods, please refer to Yao and Liu, (2025).
In this section, we apply several two-sample MR methods to perform xMR analysis, aiming to identify putative causal plasma proteins associated with the risk of Alzheimer’s disease. For the exposure, we use UK Biobank Pharma Proteomics Project (UKB‑PPP) summary statistics on 1,463 plasma proteins measured in 54,306 individuals (Sun et al., 2023b, ). For the outcome, we use summary statistics from a meta-analysis of GWASs for clinically diagnosed AD and AD by proxy, comprising 455,258 samples in total (Jansen et al.,, 2019). Genetic instruments for each protein are selected by applying a Bonferroni‐corrected threshold of $p\text{-value}<3.40\times 10^{-11}$, followed by LD clumping at threshold $r^{2}<0.01$, as described in Sun et al., 2023b . We compare the following two-sample MR methods in this application: (1) inverse-variance weighted (IVW) method (Bowden et al.,, 2017); (2) MR-Egger regression (Bowden et al.,, 2015); (3) MR using the Robust Adjusted Profile Score (MR-RAPS) (Zhao et al.,, 2020); (4) MR Pleiotropy RESidual Sum and Outlier test (MR-PRESSO) (Verbanck et al.,, 2018); (5) the weighted median method (Bowden et al.,, 2016); (6) the mode-based estimation (Hartwig et al.,, 2017); (7) MRMix (Qi and Chatterjee,, 2019); (8) the contamination mixture method (Burgess et al.,, 2020); and (9) MR with valid IV Selection and Post-selection Inference (MR-SPI) (Yao et al.,, 2024). The code for this application can be found in “DataExamples.R” within the Supplementary Files.
Proteins identified to be significantly associated with Alzheimer’s disease after Bonferroni correction (Bland and Altman,, 1995) are summarized in Figure S1 in Supplementary Materials. In Figure S1(A), we report the number of significant plasma proteins detected by each method, which ranges from 0 (MR‑PRESSO) to 14 (MRMix). MR‑PRESSO detects no significant proteins, likely because the small number of candidate IVs per protein limits its power to perform the outlier test and detect invalid instruments. In Figure S1(B), we list the 11 plasma proteins identified by at least two methods. Notably, the seven proteins identified by MR‑SPI correspond to the top seven proteins ranked by the number of supporting methods. For a more detailed discussion of these seven proteins including gene ontology analysis and AlphaFold3‑based structural prediction (Abramson et al.,, 2024), please refer to Yao et al., (2024).
@@ 12.1 Binary and survival outcomes
In epidemiological research, the binary and survival outcomes are common, yet current MR methods predominantly focus on continuous outcomes. For one-sample MR with individual-level data, establishing identification conditions for MR analysis with binary or survival outcomes is essential, particularly in the presence of invalid instruments (Clarke and Windmeijer,, 2010; Deng et al.,, 2023; Liu et al.,, 2025; Bu et al.,, 2025). For two-sample MR with summary-level data, the current standard practice is to directly apply existing two-sample MR methods under the ALICE model framework for continuous outcomes to analyze summary statistics of binary or survival outcomes, while the interpretation of the causal effect estimate obtained by this direct application is unclear and requires further justification (Zhao et al., 2019b, ).